For the following table assume we are at the end of week 7, find the cost and schedule variances. Also find CPI and SPI. Finally, calculate the ETC and EAC for the project. Activity | Predecessors | Duration | Budget. S | Actual Cost. | % Complete | (weeks) 300 600 400 450 540 340 40020 20020 100 100 100 none none 400 b, c
Answer #1

Please find below the response.

AB CDEFGHIJ Planned completion by Completed by Budget $ 2 Activity 1 23 45 6 3 A 9 Activity week 7 Week 7 100% 100% 100% 100% 100% 100% 300 6 D 600 D-80% by week 7 E-50% by week 7 2300 Solution, Calculations at week 7 PV PV at week 10 Activi Actual cost (AC| Earned val (EV) |SV-EV-PV) | SPI = EV/PV |CV = (EV-AC) | CPI= EV/AC | EAC 0.67 1.11 1.18 0.30 0.40 0.78 300 300 1.00 1.00 1.00 0.25 0.40 0.76 150 600 120 360 120 280 120 -430 16 17 18 1980 1930 There are two methods for calculating EAC, method 1 is when project has one time variance. Method 2 is used when you expect budget to continue with variance. I have used both methods as we do not know if variance will continue or if it is one time 19 2,730.00 S 2,959.33 $800.00 1 EAC actual costs (AC) +budget at completion (BAC)-earned value (EV) 21 EAC budget at completion (BAC) cost performance index (CPI) 3 ETC- estimate at completion (EAC)-actual costs (AC) $ 1,029.33 (Depends which method of EAC is considered)

Formulas